2. ∫ 12π‘₯Β² √4π‘₯Β³ + 7𝑑π‘₯




1
Expert's answer
2022-06-01T02:57:12-0400
"\int12x^2\sqrt{4x^3+7}dx"

"u=4x^3+7, du=12x^2dx"


"\int12x^2\sqrt{4x^3+7}dx=\int\sqrt{u}du=\dfrac{2}{3}u^{3/2}+C"

"=\dfrac{2}{3}(4x^3+7)^{3/2}+C"